1) Company XYZ has a monopoly on a new product. The product can be manufactured in either of two factories. The costs of production for the two factories are LaTeX: C_1=5Q^2_1C 1 = 5 Q 1 2 and LaTeX: C_2=10Q^2_2C 2 = 10 Q 2 2. The firm’s demand for the product is P = 500- 4Q. a. (10) Find the quantity produced by factory 1, the quantity produced by factory 2, the total quantity produced, and the price that maximize profit. b. (10) There is an increase in the cost of labor in Factory 1 but the cost of labor in Factory 2 remains the same. How will this increase in labor cost affect the quantity produced in Factory 1? the quantity produced in Factory 2? Total output? Price?
Answer #1

**PLEASE UPHOLD MY ANSWER, THANK YOU**

solution o- criven p= 500-4Q and Q=Qit QR Therefore total revenue is TR =PQ = (500-400Q 402 A TR = 500Q Marginal Revenue MR =Q2 = 250-90) Ч > Q2= 62-5-2-25Q, Now profit maximizing condition for 2 is MR= MC2 > 500-80,-802 = 2002 > -80,- 28Q2 = 500 20,*price change = P =500-4Q 500-4 (34-091) 500- 136-364 =$363636 = $363.636since there is an increase in cost of Cubor in fauctory 1, MC curve will move (or) shift upward Therefore, Me will cut MR, at